# Spin-triplet pair density wave superconductors

### v1

@@ Results
with $\phi_{\tilde{d}}=\arctan\left(\frac{\tilde{d}_{y}}{\tilde{d}_{x}}\right)$, which is exactly the $d$-vector for the KFF state. On the other hand, if $\mathbf{\tilde{d}}$ is along the $z$ direction corresponding to the opposite-spin pairing, $\Delta_{i,\sigma\bar{\sigma}}=\tilde{\Delta}_{i,\sigma\bar{\sigma}}e^{-i\sigma\frac{Q}{2}}=-\tilde{d}_{z}e^{-i\sigma\frac{Q}{2}}$. This corresponds to the uniform pairing order parameter with $\psi=-|\tilde{d}_{z}|\sin(\frac{Q}{2})$ and $\mathbf{d}=-i\left\{0,0,|\tilde{d}_{z}|\cos(\frac{Q}{2})\right\}$, which describes the MP state. Therefore, the finite momentum pairing state and the uniform pairing state along the vertical phase boundary are mapped onto the $p$-wave pairing state with in-plane and out-of-plane $d$-vectors through a gauge transformation, which are always degenerate in energy. This means as long as $\mu$ is close to half-filling, the vertical phase boundary always exists as shown in Fig. 3 and Fig. 5a. On the other hand, if $\mu$ is close to the band top or bottom which leads to $s$-wave paring ground state in the transformed basis with pairing order parameter $\tilde{\Delta}_{i,\sigma\sigma^{\prime}}=\left\langle d_{i\sigma}d_{i+1\sigma^{\prime}}\right\rangle=\left(\tilde{\psi}i\sigma_{y}\right)_{\sigma\sigma^{\prime}}$, the states in the original basis become $\Delta_{i,\sigma\bar{\sigma}}=\tilde{\Delta}_{i,\sigma\bar{\sigma}}e^{-i\sigma\frac{Q}{2}}=\sigma\tilde{\psi}e^{-i\sigma\frac{Q}{2}}$, corresponding to the uniform MP state with pairing order parameters described by $\psi=\tilde{\psi}\cos(\frac{Q}{2})$ and $\mathbf{d}=-i\left\{0,0,\tilde{\psi}\sin(\frac{Q}{2})\right\}$. In this case, there is no vertical phase boundary in the phase diagram, and only a small enough $r_{V}$ can drive the system into a t-PDW state. Therefore, most parts of the phase diagram are occupied by the uniform MP state with the t-PDW state shrinking to the left bottom as shown in Fig. 5b.
Discussion
In this work, we systematically study the effective 1D model with mixed SOC and nearest-neighbor attraction. We find that the triplet pairing PDW state whose $d$-vector modulates in real space can be realized in this microscopic model. Moreover, the parameter region to realize such a PDW state is large when the chemical potential $\mu$ is close to half-filling and shrinks significantly when $\mu$ approaches the band top or bottom, which is accompanied by the existence of the topological phase transition within the MP state. This work introduces a simple and concrete theoretical model to realize the PDW states and it also provides a possible explanation for the PDW state reported in the 1D domain wall of the iron-based superconductor Fe(Te,Se). We hope it stimulates the exploration of the PDW state in the embedded quantum structure of the high-temperature superconductors.
@@ Methods
The self-consistent equations can be solved in the folded Brillouin Zone (BZ) for the commensurate Q, assuming Q=$\frac{\pi}{D}$, the BZ is folded into $\frac{1}{2D}$ of the original BZ. Then the mean-field Hamiltonian should be written in the Nambu basis $\psi_{k}^{\dagger}=\left(c_{k+(n-1)Q\uparrow}^{\dagger},c_{k+(n-1)Q\downarrow}^{\dagger},c_{-k-(n-1)Q\uparrow},c_{-k-(n-1)Q\downarrow}\right)$, with $n\in[1,2D]$, as
@@
for $n\in[1,2D]$ and $\left(2n+1\right)^{\prime}=\left(2n\right)\mod(4D)+1$ and $\left(2n+2\right)^{\prime}=\left(2n+1\right)\mod(4D)+1$, so that all the indices are within the range of [1,4D].
@@
Another feature of the phase diagram when the chemical potential is tuned away from the half-filling is the existence of an extra topological phase transition within the MP state with the phase boundary represented by the red lines in Fig. 5a and Fig. 5b. These phase boundaries separate the topological superconductor and the trivial superconductor, whose topological property can be determined by a $Z_{2}$ invariant $\mathcal{N}$ [80], which was used to study the topological phase diagram in the atomic line defect [78].
We carefully study the phase transitions for these two cases. We find that for $\mu=-1.4$ where the vertical phase boundary mentioned above still exists, the topological phase transition is first-order. To see it more transparently, we fix the ratio $r_{\rm so}=1.0$ and tune the parameter $r_{V}$ through the phase boundary. We find the total energy has a hysteresis behavior as the parameter $r_{V}$ is tuned from big to small values and the small to big values respectively which is shown in Fig. 6a, indicating the first-order feature of the transition. Moreover, the band structure of the TSC and trivial SC state are almost identical close to the transition point and the band gap of the system never closes across the phase transition as shown in Fig. 6c.
Figure 6: Topological phase transitions within the MP states. a,b The evolution of the total energy across the phase boundary between the topological trivial and non-trivial MP states shown in Fig. 3b,c with the fixed value of $r_{\rm so}=1.0$ and $\mu$=-1.4 for a and $r_{\rm so}=3.0$ and $\mu$=-1.8 for b. Here, the black (red) line corresponds to tuning $r_{V}$ from big (small) to small (big) values, and the hysteresis behavior in a indicates the 1st-order character of the phase transition. c,d The evolution of the extracted band gap of the states across the corresponding phase boundary described in a,b. The band gap never closes for the case with $\mu=-1.4$ (c) and closes and reopens across the phase boundary with the critical value $r_{V}$=0.348 for the case with $\mu=-1.8$ (d).
@@ Acknowledgments
YZ is supported in part by National Natural Science Foundation of China (NSFC) Grants No. 12074276 and No. 12274279. ZW is supported by the U.S. Department of Energy, Basic Energy Sciences, Grant No. DE FG02-99ER45747.

### v2

@@ Results
with $\phi_{\tilde{d}}=\arctan\left(\frac{\tilde{d}_{y}}{\tilde{d}_{x}}\right)$, which is exactly the $d$-vector for the KFF state. On the other hand, if $\mathbf{\tilde{d}}$ is along the $z$ direction corresponding to opposite-spin pairing, $\Delta_{i,\sigma\bar{\sigma}}=\tilde{\Delta}_{i,\sigma\bar{\sigma}}e^{-i\sigma\frac{Q}{2}}=-\tilde{d}_{z}e^{-i\sigma\frac{Q}{2}}$. This corresponds to the uniform pairing order parameter with $\psi=-|\tilde{d}_{z}|\sin(\frac{Q}{2})$ and $\mathbf{d}=-i\left\{0,0,|\tilde{d}_{z}|\cos(\frac{Q}{2})\right\}$, which describes the MP state. Therefore, the finite momentum pairing state and the uniform pairing state along the vertical phase boundary are mapped onto the $p$-wave pairing state with in-plane and out-of-plane $d$-vectors through a gauge transformation, which are always degenerate in energy. This means as long as $\mu$ is close to half-filling, the vertical phase boundary always exists as shown in Fig. 3 and Fig. 5a. On the other hand, if $\mu$ is close to the band top or bottom which leads to $s$-wave paring ground state in the transformed basis with pairing order parameter $\tilde{\Delta}_{i,\sigma\sigma^{\prime}}=\left\langle d_{i\sigma}d_{i+1\sigma^{\prime}}\right\rangle=\left(\tilde{\psi}i\sigma_{y}\right)_{\sigma\sigma^{\prime}}$, the states in the original basis become $\Delta_{i,\sigma\bar{\sigma}}=\tilde{\Delta}_{i,\sigma\bar{\sigma}}e^{-i\sigma\frac{Q}{2}}=\sigma\tilde{\psi}e^{-i\sigma\frac{Q}{2}}$, corresponding to the uniform MP state with pairing order parameters described by $\psi=\tilde{\psi}\cos(\frac{Q}{2})$ and $\mathbf{d}=-i\left\{0,0,\tilde{\psi}\sin(\frac{Q}{2})\right\}$. In this case, there is no vertical phase boundary in the phase diagram, and only a small enough $r_{V}$ can drive the system into a t-PDW state. Therefore, most parts of the phase diagram are occupied by the uniform MP state with the t-PDW state shrinking to the left bottom as shown in Fig. 5b.
Conclusions
In this work, we systematically study the effective 1D model with general SOC and nearest-neighbor attraction. We find that the triplet pairing PDW state whose $d$-vector modulates in real space can be realized in this microscopic model. Moreover, the parameter region to realize such a PDW state is large when the chemical potential $\mu$ is close to half-filling and shrinks when $\mu$ approaches the band top or bottom, which is accompanied by the topological phase transition within the MP state. This work introduces a simple and concrete theoretical model to realize the PDW states. It also provides a possible explanation for the PDW state reported along the 1D domain wall in monolayer iron-based superconductor Fe(Te,Se) grown on STO. We hope it will stimulate the exploration of the PDW state in embedded quantum structures of high-temperature superconductors.
@@ Methods
[…]
@@
for $n\in[1,2D]$ and $\left(2n+1\right)^{\prime}=\left(2n\right)\mod(4D)+1$ and $\left(2n+2\right)^{\prime}=\left(2n+1\right)\mod(4D)+1$, so that all the indices are within the range of [1,4D]. The self-consistent loop begins with an initial guess for the order parameters $\left(\Delta_{\parallel,0},\Delta_{\parallel,Q},\Delta_{\parallel,-Q},\Delta_{\perp,0},\Delta_{\perp,Q},\Delta_{\perp,-Q}\right)$. Using these parameters, the meanfield Hamiltonian (Eq.15) is diagonalized in the folded BZ. The resulting ground-state expectation values are then computed using Eq.16 to obtain an updated set of order parameters. These updated parameters are linearly mixed with the previous set and used to construct the mean-field Hamiltonian for the next iteration. The process is repeated until convergence, defined as the numerical equality of the old and updated order parameters within a specified tolerance.
@@
Another feature of the phase diagram when the chemical potential is tuned away from the half-filling is the existence of an extra topological phase transition within the MP state with the phase boundary represented by the red lines in Fig. 5a and Fig. 5b. These phase boundaries separate the TSC and the trivial superconductor, whose topological property can be determined by a $Z_{2}$ invariant $\mathcal{N}$ [91], which was used to study the topological phase diagram in the atomic line defect [89].
We carefully study the phase transitions for these two cases. We find that for $\mu=-1.4$ where the vertical phase boundary mentioned above still exists, the topological phase transition is first-order. To see it more transparently, we fix the ratio $r_{\rm so}=1.0$ and tune the parameter $r_{V}$ through the phase boundary. We find the total energy has a hysteresis behavior as the parameter $r_{V}$ is tuned from big to small values and the small to big values respectively which is shown in Fig. 6a, indicating the first-order feature of the transition. Moreover, the band structure of the TSC and trivial SC state are almost identical close to the transition point and the band gap of the system never closes across the phase transition as shown in Fig. 6c. Across the topological phase transitions without band gap closure [92], the time-reversal symmetry can be spontaneously broken by mixing the degenerate topological trivial and non-trivial MP states [93]. We perform the meanfield calculation including TRSB orders and find that such solutions indeed exist. Near the topological phase transition points, where the two MP states are degenerate in energy, these TRSB states become energetically favored over the TRI ones as shown in the inset of Fig. 6a. These states mix the $s$- and $p$- wave pairing with a finite phase difference, giving rise to the ss+ipp states, as depicted by the narrow shaded grey region bounded by the red lines in Fig. 5a. Away from the degenerate points, the ss+ipp states become less favored and the ground states revert to the TRI MP states. We thus conclude that the TRSB ss+ipp states emerge in a narrow region near the phase transition and provide an intriguing pathway across the topological phase transition between two TRI MP states.
Figure 6: Topological phase transitions within the mixed-parity (MP) states. a, b The evolution of the total energy across the phase boundary between the topological trivial and non-trivial MP states with the fixed value of $r_{\rm so}=1.0$ and $\mu$=-1.4 for a and $r_{\rm so}=3.0$ and $\mu$=-1.8 for b. Here, the black (red) line corresponds to tuning $r_{V}$ from big (small) to small (big) values indicated by the black (red) arrows, showing the hysteresis behavior in a only. The green line in a represents the energy of the time-reversal symmetry-breaking (TRSB) states across the phase boundary. The inset shows a zoomed-in region around the transition points where the two MP states are degenerate, indicating that the TRSB state can become the ground state by mixing the two MP states with a nonzero relative phase near the degeneracy point. The two vertical dashed lines indicate the range where the TRSB state becomes the ground state for a fixed value of $r_{so}$. c, d The evolution of the extracted band gap of the states across the corresponding phase boundary described in a and b. The band gap never closes for the case with $\mu=-1.4$ (c) and closes and reopens across the phase boundary with the critical value $r_{V}$=0.348 for the case with $\mu=-1.8$ (d).
@@
Topological invariant for the MP and PDW states
Since the 1D model respects both the time-reversal and particle-hole symmetry, it belongs to class $DIII$ [94] whose topological property can be described by a $Z_{2}$ invariant $\mathcal{N}$ [91]. A simple expression for computing this invariant involves the expectation value of the time-reversed pairing functions for each band $n$ at momentum $k$:
$\delta_{nk}=\langle n,k|{\cal T}h_{\Delta}^{\dagger}(k)|n,k\rangle$ (29)
where $|n,k\rangle$ denotes the eigenstate of the normal state Hamiltonian $h_{N}(k)$, $\cal T$} is the time-reversal matrix satisfying ${\cal T}^{\dagger}h_{N}(k){\cal T}=h_{N}^{T}(-k)$ and $h_{\Delta}(k)$ is pairing matrix. In the weak pairing limit, the topological invariant $\cal N$} in 1D is given by the product
${\cal N}=\Pi_{s}[{\rm sgn}(\delta_{s})]$ (30)
where $s$ runs over all Fermi points of all bands between 0 and $\pi$.
For the MP state where only the two order parameters $\Delta_{\parallel,0}$ and $\Delta_{\perp,0}$ are finite, the pairing matrix becomes
$\begin{split}h_{\Delta}(k)&=V_{1}\left(\text{Re}[\Delta_{\parallel,0}]i\sigma_{0}-\text{Im}[\Delta_{\parallel,0}]\sigma_{z}\right)\sin{k}\\ &+V_{2}\left(\text{Im}[\Delta_{\perp,0}]\sigma_{x}\sin{k}-\text{Re}[\Delta_{\perp,0}]i\sigma_{y}\cos{k}\right),\end{split}$ (31)
the time-reversal matrix ${\cal T}=-i\sigma_{y}$ and $h_{N}(k)$ is $h_{0}(k)$ defined in Eq. 5. Then Eqs. 29,30 can be directly used to calculate the $Z_{2}$ invariant $\cal N$}, which leads to the topological phase transition shown in Fig. 5.
For the PDW state with $Q=\frac{\pi}{D}$, where the finite order parameters are $(\Delta_{\parallel,Q},\Delta_{\parallel,-Q},\Delta_{\perp,Q},\Delta_{\perp,-Q})$, the above method can be generalized to calculate the $Z_{2}$ invariant $\cal N$}. In order to express the time-reversal matrix more conveniently, we choose the basis as $\phi_{k}^{\dagger}=\left(c_{k-\frac{2D-1}{2}Q\uparrow}^{\dagger},c_{k-\frac{2D-1}{2}Q\downarrow}^{\dagger},c_{k-\frac{2D-3}{2}Q\uparrow}^{\dagger},c_{k-\frac{2D-3}{2}Q\downarrow}^{\dagger},\cdots,c_{k+\frac{2D-1}{2}Q\uparrow}^{\dagger},c_{k+\frac{2D-1}{2}Q\downarrow}^{\dagger}\right)$ so that the Nambu basis is defined as $\Psi_{k}^{\dagger}=\left(\phi_{k}^{\dagger},\phi_{-k}^{T}\right)$ and $k$ is defined in the folded BZ, i.e. $k\in[-\frac{Q}{2},\frac{Q}{2}]$. In this basis, the time-reversal matrix becomes ${\cal T}=J_{2D}\otimes(-i\sigma_{y})$, where $J_{2D}$ is the counteridentity matrix whose elements are all equal to zero except those on the antidiagonal, which are all equal to 1 with dimension $2D$ and the normal state Hamiltonian $h_{N}(k)$ becomes block-diagonal which can be expressed as
$h_{N}(k)=\bigoplus_{i=1}^{2D}h_{0}\left(k+\frac{2i-2D-1}{2}Q\right).$ (32)
Moreover, the pairing matrix $h_{\Delta}(k)$ becomes a matrix of size 4D, whose block form can be expressed as
$\begin{cases}h_{\Delta}(k)_{n,2D-n}=h_{p,-}\left(k+\frac{2n-2D-1}{2}Q\right),n\in[1,2D-1]\\ h_{\Delta}(k)_{n,2D+2-n}=h_{p,+}\left(k+\frac{2n-2D-1}{2}Q\right),n\in[1,2D-1]\\ h_{\Delta}(k)_{2D,2D}=h_{p,-}\left(k+\frac{2D-1}{2}Q\right)\\ h_{\Delta}(k)_{1,1}=h_{p,+}\left(k+\frac{1-2D}{2}Q\right)\end{cases}$ (33)
where,
$h_{p,-}(k)=\left[\begin{array}[]{cc}2iV_{1}\Delta_{\parallel,-Q}\sin(k+\frac{1}{2}Q)&-V_{2}\left[\Delta_{\perp,-Q}e^{i(k+\frac{1}{2}Q)}+\Delta_{\perp,Q}^{*}e^{-i(k+\frac{1}{2}Q)}\right]\\ V_{2}\left[\Delta_{\perp,-Q}e^{-i(k+\frac{1}{2}Q)}+\Delta_{\perp,Q}^{*}e^{i(k+\frac{1}{2}Q)}\right]&2iV_{1}\Delta_{\parallel,Q}^{*}\sin(k+\frac{1}{2}Q)\end{array}\right]$ (34)
$h_{p,+}(k)=\left[\begin{array}[]{cc}2iV_{1}\Delta_{\parallel,Q}\sin(k-\frac{1}{2}Q)&-V_{2}\left[\Delta_{\perp,-Q}^{*}e^{-i(k-\frac{1}{2}Q)}+\Delta_{\perp,Q}e^{i(k-\frac{1}{2}Q)}\right]\\ V_{2}\left[\Delta_{\perp,-Q}^{*}e^{i(k-\frac{1}{2}Q)}+\Delta_{\perp,Q}e^{-i(k-\frac{1}{2}Q)}\right]&2iV_{1}\Delta_{\parallel,-Q}^{*}\sin(k-\frac{1}{2}Q)\end{array}\right]$ (35)
Then the $Z_{2}$ invariant can be calculated using Eq. 29,30, with $s$ running over all Fermi points of all bands between 0 and $\frac{Q}{2}$. For the t-PDW state, the $Z_{2}$ invariant $\cal N$} is always -1 indicating that the t-PDW state is indeed a TRI TSC, which is consistent with the two pairs of MZMs found in the finite length chain. In this specific case, the eigenstates of the normal Hamiltonian are doubly degenerate, we thus introduce a small perturbation proportional to ${\cal T}h_{\Delta}^{\dagger}(k)$ to lift the degeneracy and the resulting eigenstates are used to calculate the $Z_{2}$ invariant.
Data availability
The data that support the findings of this study are available from the corresponding authors upon request.
Code availability
The source code of the mean-field calculation is available from the corresponding author on reasonable request.
@@ Acknowledgments
YZ is supported in part by the Shanghai Science and Technology Innovation Action Plan (Grant No. 24LZ1400800) and National Natural Science Foundation of China (NSFC) Grants No. 12274279. ZW is supported by the U.S. Department of Energy, Basic Energy Sciences, Grant No. DE FG02-99ER45747.
Author contributions
Y.Z. and Z.W. contributed to all aspects of this work and wrote the paper.
Competing interests
The authors declare no competing interests.